# The Initial Value Problem for Harmonic maps of Cohomogeneity One manifolds

### v2

@@ 1. Introduction
Organization: In Section 2 we provide preliminaries. The initial value problem for harmonic maps of cohomogeneity one manifolds is set up and solved in Section 4. The initial value problem for biharmonic maps of (specific) cohomogeneity one manifolds is set up and solved in Section 5. Finally, in Section 6, we study regular-singular systems of first order.
@@ 4. The initial value problem for harmonic maps
We rewrite (4.1) as system of first order equations:
$\dot{r}(t)=s(t),$
$\dot{s}(t)=-\dim\mathfrak{p}\frac{s(t)}{t}+\dim\mathfrak{p}\frac{r(t)}{t^{2}}-f_{5}(t,r(t),s(t)),$ (4.2)
with initial conditions $r(0)=0$ and $s(0)=v$ for some $v\in\mathbb{R}$. This system can also be rewritten as
$\dot{r}(t)=s(t),$
$\dot{s}(t)=\frac{f_{6}(t,r(t),s(t))}{t}+\dim\mathfrak{p}\frac{r(t)}{t^{2}},$ (4.3)
for $f_{6}:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ smooth.
Following [10] and [9], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (4) depending continuously on $v$. This method has been discussed in [9] for a large class of systems, in particular (5.1) falls into the category of systems discussed there. This machinery works since $f_{6}(t,r,s)$ is bounded and uniformly Lipschitz in the variables $r$ and $s$ under the condition that $r$ and svs-v are bounded by some sufficiently small constant, whence the claim. ∎
Remark 4.3.
  1. (1)
@@
    is a regular-singular ordinary differential equation.
@@ 5. The initial value problem for biharmonic maps
Clearly, we obtain the same ODE if the porbit space $M/G$ is isometric to $0,\infty)$. In this section we solve the initial value problem for biharmonic maps at singular orbits:
@@
Following again [10] and [9], one can use a modified Picard iteration to show that for each $v$ there exists a unique smooth solution of (5.1) depending continuously on $v$. ∎
@@ 6. Regular-singular systems of first order
We refer to [14, 23, 25, 15] and the references therein.

### v3

@@ 1. Introduction
Organization: In Section 2 we provide preliminaries. We provide the tension field for equavariant smooth maps between cohomogeneity one manifolds in Section 3. The initial value problem for harmonic maps of cohomogeneity one manifolds is set up and solved in Section 4. The initial value problem for biharmonic maps of (specific) cohomogeneity one manifolds is set up and solved in Section 5. Finally, in Section 6, we study regular-singular systems of first order.
@@ 3. Harmonic maps between cohomogeneity one manifolds
We apply the preceding results to study conformal changes of the domain metric.
Lemma 3.6.
Let $(M,g)$ be a cohomogeneity one manifold. Further, let $\psi:(M,\exp(2\alpha(t))g)\rightarrow(M,g)$ be the equivariant map given by
$\psi(g\cdot\gamma(t))=g\cdot{\gamma}(r(t)),$
where $r:(M/G)^{\circ}\rightarrow\mathbb{R}$ which in addition satisfies boundary conditions in cases (3) and (4). Then we have
$\tau^{\mathrm{nor}}_{|\gamma(t)}=\ddot{r}(t)+\tfrac{1}{2}\dot{r}(t)\operatorname{Tr}P_{t}^{-1}\dot{P}_{t}+\dim(\mathfrak{n})\dot{\alpha}(t)\dot{r}(t)-\tfrac{1}{2}\operatorname{Tr}P_{t}^{-1}(\dot{P})_{r(t)}.$
Proof.
This follows from Theorem 3.3. ∎
@@ 4. The initial value problem for harmonic maps
We will reduce the initial value problem to an initial value problem for which a solution theory exists. The following theorem will be applied:
Theorem 4.2 ([11]).
Consider the singular initial value problem
$\dot{y}=\frac{1}{t}M_{-1}(y)+M(t,y),\qquad y(0)=y_{0},$ (4.1)
where $y$ takes values in $\mathbb{R}^{k}$, $M_{-1}:\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is a smooth function of $y$ in a neighborhood of $y_{0}$ and $M:\mathbb{R}\times\mathbb{R}^{k}\rightarrow\mathbb{R}^{k}$ is smooth in $t,y$ in a neighborhood of $(0,y_{0})$. Assume that
  1. (1)
    $M_{-1}(y_{0})=0$;
  2. (2)
    $h\text{Id}-d_{y_{0}}M_{-1}$ is invertible for all $h\in\mathbb{N}$, $h\geq 1$.
Then there exists a unique solution $y(t)$ of (4.1). Furthermore $y$ depends continuously on $y_{0}$ satisfying (i) and (ii).
@@
is a regular-singular ordinary differential equation whose solutions are given by
$r(t)=c_{0}t+c_{1}t^{-\dim\mathfrak{p}},$
where $c_{0},c_{1}\in\mathbb{R}$. Clearly, the solution $r$ extends smoothly to $t=0$ if and only if $c_{1}=0$.
We thus make the Ansatz
$r(t)=c_{0}ta(t),$ (4.3)
where $a:(0,\infty)\rightarrow\mathbb{R}$ is smooth. We plug this Ansatz into (4.2) and hence we obtain
$a^{\prime\prime}(t)+\frac{p+2}{t}a^{\prime}(t)=f(t,a(t),a^{\prime}(t)),$
where $f:[0,\infty)\times\mathbb{R}^{2}\rightarrow\mathbb{R}$ is a smooth map. We let $u(t)=a^{\prime}(t)$ and thus obtain the following first order system
$\frac{d}{dt}\begin{pmatrix}a(t)\\ u(t)\end{pmatrix}=\frac{1}{t}\begin{pmatrix}0&0\\ 0&-(p+2)\end{pmatrix}\begin{pmatrix}a(t)\\ u(t)\end{pmatrix}+\begin{pmatrix}u(t)\\ f(t,a(t),u(t))\end{pmatrix}$. (4.4)
We observe that
$h\operatorname{Id}_{2}-\begin{pmatrix}0&0\\ 0&-(p+2)\end{pmatrix}$
is invertible for each $h\in\mathbb{N}=\{1,2,3\dots\}$ and let $a(0)=c_{1}$, $u(0)=0$. Hence Theorem 4.2 yields that there exists a unique solution
$y(t)=\begin{pmatrix}a(t)\\ u(t)\end{pmatrix}$
of 4.4 with
$y(0)=\begin{pmatrix}c_{1}\\ 0\end{pmatrix}$.
Furthermore $y$ depends continuously on $c_{1}$. Combining with (4.3) yields the claim. ∎
Remark 4.4.
@@ 5. The initial value problem for biharmonic maps
Clearly, we obtain the same ODE if the orbit space $M/G$ is isometric to $0,\infty)$. In this section we solve the initial value problem for biharmonic maps at singular orbits:
@@
We proceed as in the proof of Theorem 4.1 to establish the claim. ∎
@@ 6. Regular-singular systems of first order
We refer to [10, 11, 15, 16, 24, 26, 9] and the references therein.